The rate of color toner increases $33 to $35 and rate of black&white toner increases $28 to $29 if company spend $1540 in which 3/5 is color toner how much is total expenditure.
Answer by Tatiana_Stebko(1539)

1. 1540*(3/5)=924($) for color toners
2. 924/33=28 color toners
3. 1540-924=616 ($) for b/w toners
4. 616/28=22 b/w toners
5. 28*35+22*29=1618 ($) total expenditure